will you help me determine the value if g(x)= 2x2 (the last 2 is an exponent)
1) g1/2

Answer by stanbon(75887)

help me determine the value if g(x)= 2x2 (the last 2 is an exponent)
1) g1/2
---------------
g(x) = 2x^2
g(1/2) = 2(1/2)^2 = 2*(1/4) = 1/2
==========================
Cheers,
Stan H.